OpenStudy (anonymous):

PLEASE HELP ME I WILL GIVE OUT BEST ANSWER AWARD PLEASE. the cost of a taxi ride depends on the distance traveled. you paid $8.50 for a 3 mile ride while your friend paid $18.50 for an 8 mile ride. What is the equation in point slope form?

OpenStudy (anonymous):

cost depends on how far you've gone, and if cost DEPENDS then it's going to be the y-value which leaves miles as the X-value X Y 3miles $8.50 8miles $18.50 remember slope is m \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] lets have y_2 = 18.5 and x_2 = 8, can you plug in these values and the other values to find the slope?

OpenStudy (anonymous):

okay my slope was 10/5 and simplified it to 2 now what do i do ? @DemolisionWolf

OpenStudy (anonymous):

good, so now we will focus on putting it into the 'slope-intercept' form, which is y = mx+b we already know what m is 2 and we can use either (3, 8.50) or (8, 18.5) for x and y in the equation, so all that is left is to find what b is so we will do this by plugging in m, x, and y into the equation y = mx+b then solve the equation for b can you do that?

OpenStudy (anonymous):

how do i do that @DemolisionWolf

OpenStudy (anonymous):

so it would be like this.. we know that: m = 2 x = 3 y = 8.5 then we plug those numbers in the standard line equation of y = m*x + b so 8.5 = 2*3 + b make sense now? can you solve for b?

OpenStudy (anonymous):

is it 8.5= 2(3) +b -6 -6 b=2.5

OpenStudy (anonymous):

yep good job!